OpenStudy (anonymous):

Can You help me please You are a member of your local movie theater’s club. Every time you see a movie at the theater, you earn 2 advantage points. When you earn 100 points, you get a free movie pass. Currently, you have 40 advantage points. Write an equation to model the number of movies m you have to watch before you earn a free movie pass.

OpenStudy (anonymous):

Current number of points = 40 Points needed = 100 - 40 = 60 For every movie, you get 2 points. So for 60 points you need to watch 60/2 = 30 movies. Try doing it now.

OpenStudy (anonymous):

So the equation would be: \[m = \frac{ 100 - 40 }{ 2 }\]

OpenStudy (mayankdevnani):

You have 40 advantage points. You still need 60 to get a free movie pass= 100-40=60

OpenStudy (mayankdevnani):

You get 2 points per movie means: 60/2=30. You have to see 30 more movies to get a free movie pass !

OpenStudy (mayankdevnani):

ok @MathLearner!

OpenStudy (anonymous):

is pratu043's equation correct?

OpenStudy (anonymous):

Well, you can write the same equation in many ways like the one I have given above. However, I would suggest you go with this one: \[2m = 60\]

OpenStudy (anonymous):

thanks @pratu043